OpenStudy (anonymous):

Solve for x: x + 5 = 1 6 3

OpenStudy (anonymous):

3(x+5)=6 x+5=6/3 x+5=2 x=2-5 x=-3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

anytime